### v1

@@ 3.18 Symmetry metric by Brachmann & Redies
Figure 12 was obtained by plotting the mean of the left-right (L-R) symmetry metric  (Brachmann2016, ) against $\mu$ for the whole database (picture A), the unmodified images only (picture B), and the boxplots of this metric for three groups (uglified & auto-uglified, unmodified and beautified in picture C).
Figure 12: Results for the Left-Right Symmetry metric (Brachmann2016, ). (A) Scatterplot for all images. (B) Same for unmodified images. (C) Metric values for the three groups of images considered. Boxplots as in previous figures.
@@ 3.19 Mean 3D depth
We applied the 3D depth algorithm (Ranftl2021, ) and calculated its mean value to obtain an estimation of depth in each image. We resized the images so that the longer side was 150 pixels to save computer processing power.
@@ 4 Discussion
It may be argued that image aesthetics are intrinsically linked with semantics, so the MSC dataset represents only a subset of the full image spectrum with characteristics that are different from those of everyday images. From our point of view, the MSC dataset can be seen as a controlled experiment designed to isolate perceptual factors from semantic ones, at least for one broad class of images. This reductionist approach is intentional and aligns with methods in neuroscience that separate complex processes to understand individual components, like the "what" (ventral) and the "where" (dorsal) pathways (Ungerleider1994, ) or specialized brain regions like visual areas V1, V2, V3 (Felleman1991, ), etc. By minimizing semantic factors and holding them constant, we provide a foundation for models that can later introduce semantic complexity. Although our dataset may not capture the full scope of real-world aesthetics, this methodology enables a structured investigation of aesthetic decisions based on perception, paving the way for future research that integrates semantic components and deepens our understanding of image aesthetics.
@@
We lack a solid explanation for the low dynamic range and the scarcity of examples categorized as extremely ugly and extremely beautiful examples of the AVA database and other databases derived from the DPChallenge. This may have to do with a combination of psychophysical errors such as the time-order error first described by Fechner (Stevens1957, ) and contextual effects like the within-category assimilation effect (Tajfel1963, ). The former refers to the fact that the order of presentation influences how stimuli are judged (Hellstrom1985, ), and the later refers to a decrease in perceived difference that occurs when two stimuli are somehow perceived as belonging to the same category (Arielli2012, ). Since we do not know for sure how the stimuli were presented to DPChallenge participants (e.g., if the images were always presented in the same order or randomized?) and, similarly, do not know the effects of the different categories (“themes” or “challenges”), it is difficult to find a satisfactory answer. Nevertheless, by presenting our images in random order, not having “themes” or “challenges” in our crowdsourcing, and creating our own “ugly” examples, we have avoided at least these two important flaws. In consequence MSC is more balanced, has similar numbers of images in all valuation categories and has no bias towards beautiful images. Indeed, when naïve observers were presented with uglified or auto-uglified images, they tended to rate them with the lowest aesthetic valuations. Figure 5 shows that modified images are generally considered uglier than unmodified images, filling a void present in classical databases. In other words, when presented with extremely ugly examples, naïve observers rescale their decisions and assign previously low-rated images much higher ratings than otherwise—a consequence of the anchoring effect (Tversky1974, ).
@@
Another crucial aspect to consider is the effectiveness of the semantic content removal in MSC. It is difficult to dispute that eliminating elements such as man-made objects, people, animals, etc., would qualitatively result in a dataset that elicits fewer cognitive and/or emotional responses. The challenge here is to assess this affirmation quantitatively, which itself merits a future analysis. Meanwhile, we must acknowledge that we cannot entirely eliminate semantics, and some observers may have a strong attachment, preference, or emotional associations towards natural scenes and objects, as supported by research linking the viewing of natural scenes to improved health/well-being outcomes (for a review, see (Bratman2019, )). Alternatively, we had to define specific criteria to select images lacking semantic content from the extensive range of examples available on the internet. The exclusion of man-made objects, people, etc., despite being arbitrary, proved to be a reasonable approximation. Importantly, the vast majority of manipulated images stayed within the set of natural images in the sense that a naïve observer would classify them as belonging to this set.
@@ 4.1 Contrast
Artists and image professionals manipulate highlights, shadows, and lighting in relation to their surroundings to create overall effects (gloomy, moody, airy, bright, etc.). The same applies to color, allowing them to create aesthetically pleasing compositions by spatially arranging different colored elements. They exploit a property of the human visual system, which encodes its input by normalizing light intensity and processing mostly contrast patterns. Even though it is unclear how much contrast weighs in the final aesthetic decision, there is consensus on the importance of chromatic and achromatic contrast, and that it needs to be calculated in a multiscale framework for complex images (Peli1990, ). Our results support these points since the highest correlation in Table 5 occurs between multiscale contrast and $\mu$.
Multiscale contrast is also inversely related to image fragmentation (Fig. 14 shows that $C_{Lab}$ decreases almost monotonically as the image gets scrambled). This is a property of the filters used, since the positive and negative parts of large-scale DoGs tend to cancel each other in highly fragmented regions. With that in mind, we can reinterpret Figure 14 and say that naïve observers showed aesthetic preferences for unfragmented images, which is in line with psychophysical results showing that mean beauty rating decreases with crowding (Zhou2022, ).
@@ 4.2 Colorfulness
We expected colorfulness to play a strong role in aesthetic valuations, since vivid colors are commonly used in image enhancement (Zamir2021, ; Zamir2017, ) and colorless pictures of natural objects are usually deemed as “boring” or “uninteresting”. Indeed, that is the case for unmodified images, where all measures of colorfulness (CIELab gamut expansion, HVS saturation and Hasler & Suesstrunk’s measure) correlate with $\mu$ (see Table 5). These correlations are absent in the AVA database presumably masked by the semantic content or because DPChallenge users (being mostly amateur photographers) have some kind of bias in this respect. For example, saturated colors are likely to be common in human-made objects and do not have the same impact as in natural objects. The inclusion of modified images reverses the correlation, mostly because informed observers did both, reduce and increase colorfulness to decrease $\mu$. This can be clearly seen for HSV saturation in appendix Figure A6, where ugly images fall into two distinct clusters of high and low saturation. Figure A4 (appendix) shows extreme examples of how informed observers manipulated colorfulness to reduce $\mu$, suggesting perhaps a dislike of extremely unnatural colors of natural objects.
@@ 4.6 Machine-learnt features
Fig. 12 shows a clear tendency for naïve observers to give lower $\mu$ values to more symmetrical images. This is true for both L-R symmetry and U-D symmetry (see appendix Figure A9) and the literature in this respect is controversial. On the one hand, researchers have repeatedly shown symmetry to play a central role in preference and beauty judgments of visual stimuli (al-Rifaie2017, ), in particular of human faces (Rhodes2006, ; Tinio2013, ), as this preference has been associated with advantages regarding mate choice (Thornhill1997, ). Symmetry is also a strong determinant of aesthetic judgments of meaningless, abstract patterns and is often explained with reference to the efficient functioning of the human visual system (Reber2004, ). On the other hand, significant differences in the valuation of symmetry has been reported between groups of art experts and non-experts (Leder2019, ). A careful examination of the examples in appendix Figure A10, shows that these results may have been a consequence of how the neural net was trained. Images at the top contain objects to one side of the scene, which in the context of music CD covers were considered highly L-R asymmetrical. Images at the bottom have low contrast (most of their features were removed), which makes them highly symmetrical and, according to our results, ugly (low $\mu$). Also, images of extensive foliage with uninteresting features may have been classified as symmetrical and low $\mu$. A cross-correlation analysis (see below) reveals that this metric is significantly correlated to others: multiscale contrast or image fragmentation ($r=-0.39$, $p<0.001$), Focus/Blur ($r=0.44$, $p<0.001$) and $\alpha$-slope ($r=0.36$, $p<0.001$), which helps to understand these results.
The last of the machine-learnt features analyzed was the mean 3D depth metric of Ranftl et al(Ranftl2021, ). Figure A11 in the appendix shows a few examples of this metric, where high values correspond to content located close to the viewer (and on the same focal plane), and low values correspond to far away objects, like clouds. Our results show that informed observers did follow the general trend in unmodified images by increasing this value to lower $\mu$. Again, there is no correlation between this metric and $\mu$ for the AVA database.
@@ 4.7 Correlations among the different metrics
The removal of semantic-rich images is likely to reduce the generalizability of the MSC database, limiting it to study the contribution of the sensory–motor (perceptual) circuitry (Chatterjee2014, ; Redies2015, ) to aesthetic judgment. Nonetheless, we believe that using a reductionist approach, such as the one proposed, is a good and essential step towards advancing understanding when facing notably intricate and complex problems. In the future we may want to explore the contribution of the meaning–knowledge and the emotion–valuation circuitry (Chatterjee2014, ; Redies2015, ) for generating ugly imagery using a similar approach. Many questions remain. For example, what is the role of emotion in the uglification process? Will informed observers implement changes similar to those we uncovered in the presence of strong, emotion-driving, semantic content?
@@ A.5. Overall Perceived Colorfulness by Hasler & Suesstrunk
Figure A5: Results for the Overall Colorfulness metric  (Hasler2003, ). (A) Scatterplot for all images. (B) Same for unmodified images.
@@ A.6. Saturation from the HSV color model
Figure A7: Results for the Focus/Blur metric (Pertuz2022, ). (A) Scatterplot for all images. (B) Same for unmodified images.
@@ A.9. Up-Down Symmetry by Brachmann & Redies
Figure A9: Results for the Up-Down Symmetry metric  (Brachmann2016, ). (A) Scatterplot for all images. (B) Same for unmodified images.
@@ A.10. Extreme examples of poor and good L-R symmetry
Figure A11: Examples of images with very high and very low 3D depth values according to Ranftl et al (Ranftl2021, ). Top panel: images with high 3D depth values. Bottom panel: images with low 3D depth values. Note: in this metric, small values indicate larger mean depth.

### v2

@@ 3.18 Symmetry metric by Brachmann & Redies
Figure 12 was obtained by plotting the mean of the left-right (L-R) symmetry metric [97] against $\mu$ for the whole database (picture A), the unmodified images only (picture B), and the boxplots of this metric for three groups (uglified & auto-uglified, unmodified and beautified in picture C).
Figure 12: Results for the Left-Right Symmetry metric [97]. (A) Scatterplot for all images. (B) Same for unmodified images. (C) Metric values for the three groups of images considered. Boxplots as in previous figures.
@@ 3.19 Mean 3D depth
We applied the 3D depth algorithm [100] and calculated its mean value to obtain an estimation of depth in each image. We resized the images so that the longer side was 150 pixels to save computer processing power.
@@ 4 Discussion
It may be argued that image aesthetics are intrinsically linked with semantics, so the MSC dataset represents only a subset of the full image spectrum with characteristics that are different from those of everyday images. From our point of view, the MSC dataset can be seen as a controlled experiment designed to isolate perceptual factors from semantic ones, at least for one broad class of images. This reductionist approach is intentional and aligns with methods in neuroscience that separate complex processes to understand individual components, like the "what" (ventral) and the "where" (dorsal) pathways [103] or specialized brain regions like visual areas V1, V2, V3 [104], etc. By minimizing semantic factors and holding them constant, we provide a foundation for models that can later introduce semantic complexity. Although our dataset may not capture the full scope of real-world aesthetics, this methodology enables a structured investigation of aesthetic decisions based on perception, paving the way for future research that integrates semantic components and deepens our understanding of image aesthetics.
@@
We lack a solid explanation for the low dynamic range and the scarcity of examples categorized as extremely ugly and extremely beautiful examples of the AVA database and other databases derived from the DPChallenge. This may have to do with a combination of psychophysical errors such as the time-order error first described by Fechner [105] and contextual effects like the within-category assimilation effect [106]. The former refers to the fact that the order of presentation influences how stimuli are judged [107], and the later refers to a decrease in perceived difference that occurs when two stimuli are somehow perceived as belonging to the same category [108]. Since we do not know for sure how the stimuli were presented to DPChallenge participants (e.g., if the images were always presented in the same order or randomized?) and, similarly, do not know the effects of the different categories (“themes” or “challenges”), it is difficult to find a satisfactory answer. Nevertheless, by presenting our images in random order, not having “themes” or “challenges” in our crowdsourcing, and creating our own “ugly” examples, we have avoided at least these two important flaws. In consequence MSC is more balanced, has similar numbers of images in all valuation categories and has no bias towards beautiful images. Indeed, when naïve observers were presented with uglified or auto-uglified images, they tended to rate them with the lowest aesthetic valuations. Figure 5 shows that modified images are generally considered uglier than unmodified images, filling a void present in classical databases. In other words, when presented with extremely ugly examples, naïve observers rescale their decisions and assign previously low-rated images much higher ratings than otherwise—a consequence of the anchoring effect [109].
@@
Another crucial aspect to consider is the effectiveness of the semantic content removal in MSC. It is difficult to dispute that eliminating elements such as man-made objects, people, animals, etc., would qualitatively result in a dataset that elicits fewer cognitive and/or emotional responses. The challenge here is to assess this affirmation quantitatively, which itself merits a future analysis. Meanwhile, we must acknowledge that we cannot entirely eliminate semantics, and some observers may have a strong attachment, preference, or emotional associations towards natural scenes and objects, as supported by research linking the viewing of natural scenes to improved health/well-being outcomes (for a review, see [110]). Alternatively, we had to define specific criteria to select images lacking semantic content from the extensive range of examples available on the internet. The exclusion of man-made objects, people, etc., despite being arbitrary, proved to be a reasonable approximation. Importantly, the vast majority of manipulated images stayed within the set of natural images in the sense that a naïve observer would classify them as belonging to this set.
@@ 4.1 Contrast
Artists and image professionals manipulate highlights, shadows, and lighting in relation to their surroundings to create overall effects (gloomy, moody, airy, bright, etc.). The same applies to color, allowing them to create aesthetically pleasing compositions by spatially arranging different colored elements. They exploit a property of the human visual system, which encodes its input by normalizing light intensity and processing mostly contrast patterns. Even though it is unclear how much contrast weighs in the final aesthetic decision, there is consensus on the importance of chromatic and achromatic contrast, and that it needs to be calculated in a multiscale framework for complex images [111]. Our results support these points since the highest correlation in Table 5 occurs between multiscale contrast and $\mu$.
Multiscale contrast is also inversely related to image fragmentation (Fig. 14 shows that $C_{Lab}$ decreases almost monotonically as the image gets scrambled). This is a property of the filters used, since the positive and negative parts of large-scale DoGs tend to cancel each other in highly fragmented regions. With that in mind, we can reinterpret Figure 14 and say that naïve observers showed aesthetic preferences for unfragmented images, which is in line with psychophysical results showing that mean beauty rating decreases with crowding [112].
@@ 4.2 Colorfulness
We expected colorfulness to play a strong role in aesthetic valuations, since vivid colors are commonly used in image enhancement [113, 114] and colorless pictures of natural objects are usually deemed as “boring” or “uninteresting”. Indeed, that is the case for unmodified images, where all measures of colorfulness (CIELab gamut expansion, HVS saturation and Hasler & Suesstrunk’s measure) correlate with $\mu$ (see Table 5). These correlations are absent in the AVA database presumably masked by the semantic content or because DPChallenge users (being mostly amateur photographers) have some kind of bias in this respect. For example, saturated colors are likely to be common in human-made objects and do not have the same impact as in natural objects. The inclusion of modified images reverses the correlation, mostly because informed observers did both, reduce and increase colorfulness to decrease $\mu$. This can be clearly seen for HSV saturation in appendix Figure A6, where ugly images fall into two distinct clusters of high and low saturation. Figure A4 (appendix) shows extreme examples of how informed observers manipulated colorfulness to reduce $\mu$, suggesting perhaps a dislike of extremely unnatural colors of natural objects.
@@ 4.6 Machine-learnt features
Fig. 12 shows a clear tendency for naïve observers to give lower $\mu$ values to more symmetrical images. This is true for both L-R symmetry and U-D symmetry (see appendix Figure A9) and the literature in this respect is controversial. On the one hand, researchers have repeatedly shown symmetry to play a central role in preference and beauty judgments of visual stimuli [115], in particular of human faces [16, 116], as this preference has been associated with advantages regarding mate choice [117]. Symmetry is also a strong determinant of aesthetic judgments of meaningless, abstract patterns and is often explained with reference to the efficient functioning of the human visual system [118]. On the other hand, significant differences in the valuation of symmetry has been reported between groups of art experts and non-experts [119]. A careful examination of the examples in appendix Figure A10, shows that these results may have been a consequence of how the neural net was trained. Images at the top contain objects to one side of the scene, which in the context of music CD covers were considered highly L-R asymmetrical. Images at the bottom have low contrast (most of their features were removed), which makes them highly symmetrical and, according to our results, ugly (low $\mu$). Also, images of extensive foliage with uninteresting features may have been classified as symmetrical and low $\mu$. A cross-correlation analysis (see below) reveals that this metric is significantly correlated to others: multiscale contrast or image fragmentation ($r=-0.39$, $p<0.001$), Focus/Blur ($r=0.44$, $p<0.001$) and $\alpha$-slope ($r=0.36$, $p<0.001$), which helps to understand these results.
The last of the machine-learnt features analyzed was the mean 3D depth metric of Ranftl et al[100]. Figure A11 in the appendix shows a few examples of this metric, where high values correspond to content located close to the viewer (and on the same focal plane), and low values correspond to far away objects, like clouds. Our results show that informed observers did follow the general trend in unmodified images by increasing this value to lower $\mu$. Again, there is no correlation between this metric and $\mu$ for the AVA database.
@@ 4.7 Correlations among the different metrics
The removal of semantic-rich images is likely to reduce the generalizability of the MSC database, limiting it to study the contribution of the sensory–motor (perceptual) circuitry [120, 73] to aesthetic judgment. Nonetheless, we believe that using a reductionist approach, such as the one proposed, is a good and essential step towards advancing understanding when facing notably intricate and complex problems. In the future we may want to explore the contribution of the meaning–knowledge and the emotion–valuation circuitry [120, 73] for generating ugly imagery using a similar approach. Many questions remain. For example, what is the role of emotion in the uglification process? Will informed observers implement changes similar to those we uncovered in the presence of strong, emotion-driving, semantic content?
@@ A.5. Overall Perceived Colorfulness by Hasler & Suesstrunk
Figure A5: Results for the Overall Colorfulness metric [84]. (A) Scatterplot for all images. (B) Same for unmodified images.
@@ A.6. Saturation from the HSV color model
Figure A7: Results for the Focus/Blur metric [88]. (A) Scatterplot for all images. (B) Same for unmodified images.
@@ A.9. Up-Down Symmetry by Brachmann & Redies
Figure A9: Results for the Up-Down Symmetry metric [97]. (A) Scatterplot for all images. (B) Same for unmodified images.
@@ A.10. Extreme examples of poor and good L-R symmetry
Figure A11: Examples of images with very high and very low 3D depth values according to Ranftl et al [100]. Top panel: images with high 3D depth values. Bottom panel: images with low 3D depth values. Note: in this metric, small values indicate larger mean depth.